5 times a number, increased by 1, is 36. find the number?
I am not good with these can you help

Answer by ptaylor(2198)


Let x=the number
Then 5x= five times the number
And 5x+1=five times the number increased by 1 and we are told that this equals 36
So:
5x+1=36 subtract 1 from both sides
5x+1-1=36-1 collect like terms
5x=35 divide both sides by 5
x=7
Ck
5*7+1=36
36=36
Hope this helps--ptaylor